OpenStudy (anonymous):

3+ log base 5 (-5x)=4

OpenStudy (anonymous):

First add 3 to both sides

OpenStudy (anonymous):

3+log base5 (-5x)=7 ?

OpenStudy (anonymous):

yep but the three is gone now so you have log_5(-5x)=7 apply base 5 exponential to both sides... \[5^{\log _{5}-5x}=5^7\] log and base 5 cancels -5x=5^7 should be easy now

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome